# TEXasGAN: Tactile Texture Exploration and Synthesis System Using Generative Adversarial Network

### v2

@@ Abstract
To create more realistic experiences in human-virtual object interactions, texture rendering has become a research hotspot in recent years. Different frequency components of designed vibrations can activate texture-related sensations due to similar receptors. However, designing specific vibrations for numerous real-world materials is impractical. Therefore, this study proposes a human-in-the-loop vibration generation model based on user preferences. To enable users to easily control the generation of vibration samples with large parameter spaces, we introduce an optimization model based on Differential Subspace Search (DSS) and Generative Adversarial Network (GAN). With DSS, users can use a one-dimensional slider to easily modify the high-dimensional latent space so that the GAN can generate desired vibrations. We trained the generative model using a open dataset of tactile vibration data and selected five types of vibrations as target samples for the generation experiment. Extensive user experiments were conducted using the generated and real samples. The results indicate that our system can generate distinguishable samples that match the target characteristics. Moreover, the results also reveal a correlation between subjects’ ability to distinguish real samples and their ability to distinguish generated samples.
@@ I Introduction
Nowadays, the virtual realm built by information technologies has become an integral part of daily life. Among various interaction technologies, haptics, which enables natural and intuitive tactile interaction with virtual objects and surfaces, can make users to engage with this information-rich virtual world and improve the user experience [1, 2]. In the research field of haptics, the rendering of texture of objects has been a hot and an active topic to augment the user experience in Virtual Reality (VR) or other application scenarios [3, 4, 5]. We can make the users feel the sensation of some certain objects or textures using vibration stimuli. Nonetheless, the precise replication of tactile textures that closely approximate real-world surfaces remains a formidable undertaking.
@@
One way to present rich tactile information is through vibration stimuli due to related tactile receptors [6]. There are two methods for generating vibration stimuli: recording the actual vibration using a sensor or microphone and playing it back, or manually designing the vibration waveform. The latter is especially useful when direct measurement is not possible, such as in video games or when enhancing specific tactile sensations for richer feedback. In this study, we focus on generating tactile sensations by tuning parameters in vibrotactile stimuli.
@@
The difficulty in designing such a vibration manually is that there are many parameters to tune [3]. The diversity and abundance of tactile textures in the physical world makes it impractical to create specific models for each texture. Therefore, developing a system that can dynamically accommodate a wide range of vibration patterns while optimizing the associated parameters is highly valuable.
@@
There have been many studies concentrated on the reconstruction of vibration signals through the utilization of spectrograms derived from sensor data capturing physical surface interactions, and also gave the optimized spectrograms that match the desired texture pattern well using generative models [4, 7]. In pursuit of a broader and more versatile spectrogram of vibrotactile signals, deep learning based generative models, particularly Generative Adversatial Network (GAN), have been enlisted. These methods can effectively optimize the objective function and yield satisfactory generative results. Building upon this, we also recognize that human expert knowledge may be another crucial factor influencing the optimization, particularly in tasks related to subjective perception.
The assistance of human knowledge allows us to treat vibration patterns that are similar in terms of human perception, even if they are different in terms of power spectrum images. Consequently, integrating human judgment and sensibility into the optimization process, a concept known as “human-in-the-loop”, becomes an appealing approach to yield results that resonate with human subjective perception, as posited in [7, 8, 9].
@@
However, optimizing vibrotactile signals, which often involves multiple parameters, presents a significant challenge for manual adjustment within the latent space. Techniques like SLS [8] enable users to manipulate multiple parameters through a one-dimensional (1-D) slider. Additionally, other interactive methods have been used to involve humans in the optimization process, such as offering different generation results for users to choose from, thereby influencing the optimization direction and generating novel textures beyond the known training dataset [7]. Considering these complexities, conventional Bayesian optimization techniques like SLS [9] may face limitations. Our goal is to provide users with a continuous exploration approach within the latent space, ensuring the optimization process is seamless and free from potential discontinuities.
@@
In this paper, we introduce TEXasGAN, a user-friendly system integrating GAN and DSS for user-driven optimization of vibrotactile stimuli, as shown in Fig. 1. A simplified 1-D slider interface was developed for easy exploration of high-dimensional tactile latent space. We also emphasized the role of human perception in optimizing and evaluating tactile feedback, highlighting the significance of integrating subjective preferences into the design process. The experiment results show that subjects achieved a 58% accuracy rate in distinguishing the categories of generated vibrations in a five-class task, compared to 74% for real vibrations. We also used a linear regression model to demonstrate a significant relationship between these two groups of accuracy rates, indicating our system’s capability to recreate distinct vibration characteristics based on user preferences.
@@
We have introduced a novel idea for optimizing tactile parameters, enabling control over high-dimensional feature spaces. By integrating various network architectures, we have created a generative model tailored specifically for optimizing tactile parameters. Despite the extensive research on human-in-the-loop approaches in various interaction methods, their application in haptics remains underexplored. We are also trying to bring We hope our work contributes to the development of more intuitive and realistic haptic feedback systems.
@@ II-A Vibration Tactile Presentation
Significant progress in vibrotactile haptic devices has made them a key focus in tactile presentation systems. Many studies aim to replicate various sensory experiences. For example, piezo actuators are used to simulate physical button sensations on touchscreens, as demonstrated by Sadia et al. [10]. Additionally, wearable devices like FingerX [11] enhance the sense of object shapes in VR environments.
Among the array of sensory modalities, the recreation of tactile textures has garnered significant attention. Through the capture of sound recordings while traversing real surfaces, the corresponding texture-specific vibrations can be faithfully reproduced via audio signals [12]. Nevertheless, it is noteworthy that this research underscores the necessity of meeting specific requirements in terms of both time resolution and frequency resolution to effectively replicate tactile sensations. Furthermore, it highlights the critical role of parameter design in shaping the characteristics of the vibrotactile signal.
@@ II-B Generative Neural Network
In addressing parameter design challenges in the domain of vibrotactile texture reproduction, a variety of optimization algorithms have been applied, as highlighted in recent works [13, 14, 15]. However, for scenarios involving a substantial number of parameters, deep learning techniques, which have seen significant advancements in recent years, have emerged as a more effective solution for managing the intricate optimization aspects of texture generation.
Autoencoders, due to their straightforward architecture and unsupervised training approach, have been used in some studies for compressing and reconstructing tactile signals [16, 17]. Nevertheless, the compression process can introduce oversmoothing, resulting in the loss of fine-grained details, making autoencoders difficult to reproduce detailed texture related vibration signals. Although increasing the dimension of the compressed latent vector can solve the problem, it also makes operations like optimization more difficult. Therefore, the use of generative models, rather than a decoder within the autoencoder framework, to reconstruct vibration signals has become a prudent choice.
Generative Adversarial Networks (GANs), a notable generative model, are widely recognized for their success in image generation and are increasingly applied to tactile reconstruction tasks [7, 4, 18]. A GAN consists of two main components: the generator and the discriminator. The generator produces samples from a specified distribution, often Gaussian noise, while the discriminator’s role is to distinguish between generated samples and real ones. Through iterative training, the discriminator’s ability to differentiate between genuine and generated samples improves, which in turn enhances the generator’s capability to produce increasingly realistic samples.
The remarkable generative capabilities of GAN enable it not only to generate samples from random noise but also from specific distributions, affording controlled sample generation opportunities. For example, prior research has explored the generation of surface images from vibrations [19] and, conversely, the generation of vibration signals from texture images [4, 18]. This characteristic of GAN, which encompasses not just reconstruction but also the generation of “new” samples conforming to a particular distribution while introducing variations, positions GAN as particularly well-suited for optimization problems in which we seek optimal solutions through continuous exploration of the latent space.
@@ II-C Human-in-the-loop Optimization
While GANs excel at generating samples that mathematically resemble real data, in some human evaluation tasks they do not inherently account for subjective cognition. In other words, does a small loss value equate to a similar perception or cognition? Consequently, many studies focused on optimization models incorporating a human-in-the-loop mode, which integrates human involvement into the iterative optimization process. This approach utilizes human expertise to guide the optimization direction.
Some human-in-the-loop systems focus solely on direct parameter optimization [20] of several parameter dimension, while other researches have explored optimization algorithms like Bayesian optimization [8, 21] to enable simultaneous control of multiple parameters. Furthermore, to overcome challenges when dealing with relatively large latent spaces, there have been some studies applying deep learning models in the optimization process to increase the capability of parameter dimension [7, 9]. Another significant challenge in the optimization of deep learning models arises from the difficulty of incorporating human decision-making throughout a large number of training epochs [7].
@@ III Methods
In this section, we will introduce the design and the implementation of our tactile vibration generation system. The whole human-in-the-loop optimization process is shown in Fig. 1. Users can change the vibration and obtain the closest result to the target by changing the slider position. First we will introduce the framework of the optimization method and then we will introduce the design, structure and training of our generative model (as shown in Fig. 2).
@@ III-A Human-in-the-loop Differential Subspace Search
To enable individuals to effectively leverage their knowledge for optimization purposes, it is imperative to ensure that users possess the ability to freely manipulate the resultant oscillatory patterns during each iteration of the optimization loop. However, exerting precise control over a multitude of parameters within a high-dimensional latent space is an intricate challenge.
@@
Here, the generator function $G$ takes an input latent vector $\bm{z}$ from the high-dimensional latent space $\bm{Z}$, and the evaluation function $E$ quantifies the quality of the generated output. In our specific context, the choice of the function $E$ is contingent upon the subjective perception of the user. The objective of the optimization is to discover the optimal latent vector $\bm{z}^{*}$ that maximizes the value of the evaluation function $E$ through human-guided exploration within the latent space [9].
@@ III-B Generator Design
We designed our vibrotactile generator architecture based on the framework of the super-resolution residual network (SRResNet) [22, 23]. Our motivation was to effectively address high-frequency image details. There is an input convolutional layer, 16 residual blocks, a mid convolutional layer, a 4-times upscale module consisting of two convolutional layers and two pixel shuffle layer, and an output convolutional layer in the SRResNet. We introduced a Fully Connected (FC) layer to the input end of the network as the resize layer for dimension transformation, so that our latent vector can be received by SRResNet. Additionally, we made an adjustment to the final convolutional layer of the network to tailor the output to the required size, specifically to generate the desired spectrogram.
@@
To ensure the effectiveness of the generator, it is essential to employ an appropriate discriminator. A discriminator that performs exceptionally well or very poorly can hinder the training of the generator. In our study, we utilized a discriminator with a structure similar to the one presented in [22]. This discriminator comprises eight convolutional layers, albeit with adjustments to the channel numbers to align it with the specific requirements of our task. The comprehensive architecture of the Generative Adversarial Network (GAN) is depicted in Fig. 3.
The generator and the discriminator was trained simultaneously, and in the training process, the generator $G$ will learn how to generate new samples closer to the real sample, while the discriminator $D$ will discriminate the generated samples. The goal of the training can be represented as follows:
@@
where $z$ means a random noise from the distribution $p(z)$ (Gaussian distribution in this research), and the $p_{data}(x)$ denotes the distribution of the real data $x$. This competitive process encourages the generator to generate samples that are indistinguishable from real data.
@@ III-C Generation with Conditional Knowledge
While GAN has outstanding generation performance, the training process is not very stable. Additionally, GAN generate samples from random noise, resulting in uncontrollable outcomes. However, our objective is to exert control over the latent space, enabling the model to generate spectrograms that closely resemble real samples belonging to specific classes. To address these challenges, we built our model combining the advantages of GANs and autoencoders, and utilized conditional information. The inspiration for our model draws from previous works such as Auxiliary Classifier GAN (ACGAN) [24], Autoencoding GAN (AEGAN) [25], and Conditional Adversarial Autoencoder (CAAE) [26] to enhance the training process and enable the model to generate samples belonging to given classes.
Our model shown in Fig. 2 contains four parts: encoder, generator, spectrogram discriminator, and latent discriminator. The encoder extracts the feature vector from the original spectrogram, while the generator utilizes this feature vector to reconstruct the spectrogram. Similar to the Adversarial Autoencoder (AAE) paradigm, a latent discriminator is employed to encourage the encoder’s output to conform to a predetermined prior distribution. This ensures that the latent space is evenly distributed according to the specified prior distribution, allowing the generator to effectively learn a mapping from this prior distribution to the spectrogram distribution, as proposed in [27]. Additionally, a determined distribution will also benefit the following optimization of DSS. To make the feature containing richer information, we also utilized the spectrogram discriminator as a classifier and applied the auxiliary classification loss, such as the ACGAN. It will work as the discriminator and classifier when receives the encoded feature, and only as the discriminator when receives the real gaussian noise.
To avoid data loss in compression and over-smoothing reconstruction, we used a generator together with the decoder loss, and use a weight value to balance them. The generator and spectrogram discriminator were structured according to the principles of CAAE. Notably, with the label as the conditional information, the generator can generate samples belonging to certain classes. And the discriminator can learn the relationship in the pairs of the spectrogram and the corresponding class, help the generator to give correct spectrograms. The training goal can be described as follows:
@@
where $D_{z}$ represents the latent discriminator, $E$ and $G$ represent the encoder and generator, $D_{s}$ represents the spectrogram discriminator, and $C$ represent the auxiliary latent classifier, respectively. $x$ and $y$ represent the real data and the label. $TV$ represents the total variation loss. $\lambda$ and $\gamma$ are the weights for the decoder loss and the total variation loss, which can balance the smoothness and high resolution [26]. In this paper, the weight setting is $\lambda=100$ and $\gamma=10$.
We initialized the encoder using a pre-trained ResNet-50 model that was trained on the ImageNet1K_V2 dataset, and introduced an FC layer to resize the spectrogram to the shape corresponding to the pre-trained model input. The generator and the spectrogram discriminator follow the network structures illustrated in Fig. 3, maintaining consistency with those architectures. The latent discriminator contains two FC layers, and 1 output layer for the discrimination and one output layer for the classification. To apply the label as conditional information, we used one-hot labels and used an FC layer to resize the label vector to the dimension corresponding to the feature map. And then the label information will be added to the feature map as a channel.

### v4

# TexSenseGAN: A User-Guided System for Optimizing Texture-Related Vibrotactile Feedback Using Generative Adversarial Network

@@ Abstract
Texture rendering has attracted significant attention as a means of creating realistic experiences in human-virtual object interactions. But in practical applications, many limited device conditions do not support the complete reproduction of spatial and temporal tactile stimuli. Different frequency components of designed vibrations can activate texture-related sensations owing to similar receptors. Therefore, we can utilize corresponding vibration signals to provide tactile feedback within the constraints of limited device environments. However, designing specific vibrations for numerous real-world materials is impractical. This study proposes a human-in-the-loop vibration generation model based on user preferences. To enable users to easily control the generation of vibration samples with large parameter spaces, we introduced an optimization model based on Differential Subspace Search (DSS) and Generative Adversarial Network (GAN). With DSS, users can employ a one-dimensional slider to easily modify the high-dimensional latent space to ensure that the GAN can generate desired vibrations. We trained the generative model using an open dataset of tactile vibration data and selected five types of vibrations as target samples for the generation experiment. Extensive user experiments were conducted using the generated and real samples. The results indicated that our system could generate distinguishable samples that matched the target characteristics. Moreover, we established a correlation between subjects’ ability to distinguish real samples and their ability to distinguish generated samples.
@@ I Introduction
Virtual realm built by information technologies has become an integral part of our daily life. Among various interaction technologies, haptics, which enables natural and intuitive tactile interaction with virtual objects and surfaces, allows users to engage with this information-rich virtual world and improves the user experience [1, 2]. In the research field of haptics, the rendering of texture of objects has been an important topic to augment the user experience in virtual reality (VR) or other application scenarios [3, 4, 5]. Texture stimuli are spatial and temporal correlated, and many studies employ parameters such as velocities, directions, and forces to establish auto-regressive models for texture modeling [6]. But there are certain requirements for this kind of presentation methods for the corresponding presentation devices, making it challenging to deploy these models on devices that are currently widespread, such as smartphones and game controllers.
@@
Rich tactile information can be presented through vibration stimuli owing to related tactile receptors [7]. This means that even if limited conditions do not allow us to create “texture” stimuli, we can still use limited devices to generate vibration signals as “texture-related” cues. There are two methods for generating vibration stimuli: recording the actual vibration using a sensor or microphone and playing it back, or manually designing the vibration waveform. The latter is useful when direct measurement is unattainable, such as in video games or when enhancing specific tactile sensations for richer feedback. In this study, we focused on generating tactile sensations by tuning parameters in vibrotactile stimuli.
@@
The difficulty in designing such a vibration manually is that there are several parameters to tune [3]. The diversity and abundance of tactile textures in the physical world makes it impractical to create specific models for each texture. Therefore, developing a system that can dynamically accommodate a wide range of vibration patterns while optimizing the associated parameters is highly valuable.
@@
Several studies have been conducted on the reconstruction of vibration signals through the utilization of spectrograms derived from sensor data capturing physical surface interactions. These studies provide the optimized spectrograms that match the desired texture pattern well using generative models [4, 8]. In pursuit of a broader and more versatile spectrogram of vibrotactile signals, deep learning based generative models, particularly Generative Adversatial Network (GAN), have been enlisted. These methods can effectively optimize the objective function and yield satisfactory generative results.
We also recognize that human expert knowledge may be another crucial factor influencing the optimization, particularly in tasks related to subjective perception. Since human subjective perception may not always align with certain mathematical statistical metrics, it can sometimes be more appropriate for the user to guide the model’s generative direction for tasks related to subjective perception. Consequently, integrating human judgment and sensibility into the optimization process, a concept known as “human-in-the-loop”, becomes an appealing approach to yield results that resonate with human subjective perception, as posited in [8, 9, 10].
@@
However, optimizing vibrotactile signals, which involves multiple parameters, presents a significant challenge for manual adjustment within the latent space. Techniques, such as Sequential Line Search (SLS) [9], enable users to manipulate multiple parameters through a one-dimensional (1-D) slider. Additionally, other interactive methods have been used to involve humans in the optimization process, such as offering different generation results for users to choose from, thereby influencing the optimization direction and generating novel textures beyond the known training dataset [8]. Considering these complexities, conventional Bayesian optimization techniques, such as SLS [9], may face limitations. Our goal is to provide users with a continuous exploration approach within the latent space, ensuring the optimization process is seamless and free from potential discontinuities.
@@
In this study, we introduced TexSenseGAN, a user-friendly system integrating GAN and DSS for user-driven optimization, focusing on the subjective “texture-related sense” guided generative model using vibration stimuli, as shown in Fig. 1. A simplified 1-D slider interface was developed for easy exploration of high-dimensional tactile latent space. We emphasized the role of human perception in optimizing and evaluating tactile feedback, highlighting the significance of integrating subjective preferences into the design process. The experiment results show that subjects achieved a 58% accuracy rate in distinguishing the categories of generated vibrations in a five-class task, compared to 74% for real vibrations. We used a linear regression model to demonstrate a significant relationship between these two groups of accuracy rates, indicating our system’s capability to recreate distinct vibration characteristics based on user preferences.
@@
We proposed a technique for optimizing tactile parameters, enabling control over high-dimensional feature spaces. By integrating various network architectures, we created a generative model tailored specifically for optimizing tactile parameters. Despite the extensive research on human-in-the-loop approaches in various interaction methods, their application in haptics remains underexplored. We believe this work will contribute to the development of more intuitive and realistic haptic feedback systems.
@@ II-A Vibration Tactile Presentation
Vibrotactile haptic devices have become a key focus in tactile presentation systems. Several studies aim to replicate various sensory experiences. For example, piezo actuators are used to simulate physical button sensations on touchscreens, as demonstrated by Sadia et al. [11]. Additionally, wearable devices, such as FingerX [12], enhance the sense of object shapes in VR environments.
Among the array of sensory modalities, the recreation of tactile textures has garnered significant attention. Through the capture of sound recordings while traversing real surfaces, the corresponding texture-specific vibrations can be reproduced via audio signals [13]. Nevertheless, note that this research underscores the necessity of meeting specific requirements in terms of both time resolution and frequency resolution to effectively replicate tactile sensations. Furthermore, it highlights the critical role of parameter design in shaping the characteristics of the vibrotactile signal.
@@ II-B Generative Neural Network
In addressing parameter design challenges in the domain of vibrotactile texture reproduction, a variety of optimization algorithms have been applied, as highlighted in recent works [14, 15]. However, for scenarios involving a substantial number of parameters, deep learning techniques, which recently have seen significant advancements, have emerged as a more effective solution for managing the intricate optimization aspects of texture generation.
Autoencoders have been used in some studies for compressing and reconstructing tactile signals owing to their straightforward architecture and unsupervised training approach [16, 17, 18]. Nevertheless, the compression process can introduce oversmoothing, resulting in the loss of fine-grained details, making autoencoders difficult to reproduce detailed texture related vibration signals. Although increasing the dimension of the compressed latent vector can solve the problem, it makes operations, such as optimization, difficult. Therefore, the use of generative models, rather than a decoder within the autoencoder framework, to reconstruct vibration signals has become a prudent choice.
Generative Adversarial Networks (GANs), a notable generative model, are widely recognized for their success in image generation and are increasingly applied to tactile reconstruction tasks [8, 4, 19]. A GAN consists of two main components: the generator and discriminator. The generator produces samples from a specified distribution, often Gaussian noise, while the discriminator’s role is to distinguish between generated samples and real ones. Through iterative training, the discriminator’s ability to differentiate between genuine and generated samples improves, which in turn enhances the generator’s capability to produce increasingly realistic samples.
The remarkable generative capabilities of GAN enable it to generate samples from random noise as well as from specific distributions, affording controlled sample generation opportunities. For example, prior research has explored the generation of surface images from vibrations [20], and conversely, the generation of vibration signals from texture images [4, 19]. This characteristic of GAN, which encompasses reconstruction as well as the generation of “new” samples conforming to a particular distribution while introducing variations, positions GAN as particularly well-suited for optimization problems in which we seek optimal solutions through continuous exploration of the latent space.
@@ II-C Human-in-the-loop Optimization
While GANs excel at generating samples that mathematically resemble real data, in some human evaluation conditions they do not inherently account for subjective cognition. In particular, we need to determine whether a small loss value equates to a similar perception or cognition. Basic computational metrics can highlight statistical differences between generated and real samples but fail to capture human subjective perception of quality. Research has been conducted to improve GAN architectures and use better loss functions to track training progress, reducing the need for developers to stare at generated samples to figure out failure modes [21]. However, the study acknowledges that they do not provide a full quantitative evaluation of generative models. More broadly, the challenge of quantitatively evaluating GAN performance is still an open problem, with no standard metrics agreed upon [22]. This highlights a gap between statistical measures of ”realness” and human subjective judgment, especially in tasks like image and audio generation where human perception still plays a key role in evaluation. Consequently, various studies have focused on optimization models incorporating a human-in-the-loop mode, which integrates human involvement into the iterative optimization process. This approach utilizes human expertise to guide the optimization direction.
Some human-in-the-loop systems focus solely on direct parameter optimization [23] of several parameter dimension, while other explored optimization algorithms, such as Bayesian optimization [9, 15], to enable simultaneous control of multiple parameters. Furthermore, to overcome challenges when dealing with relatively large latent spaces, there have been some studies applying deep learning models in the optimization process to increase the capability of parameter dimension [8, 10]. Another significant challenge in the optimization of deep learning models originates from the difficulty of incorporating human decision-making throughout several training epochs [8].
@@ III Methods
In this section, we will introduce the design and the implementation of our tactile vibration generation system. The whole human-in-the-loop optimization process is shown in Fig. 1. Users can change the vibration and obtain the closest result to the target by changing the slider position. First, we will introduce the framework of the optimization method, and then introduce the design, structure and training of our generative model (as shown in Fig. 2).
@@ III-A Human-in-the-loop Differential Subspace Search
To enable individuals to effectively leverage their knowledge for optimization purposes, it is important to ensure that users possess the ability to freely manipulate the resultant oscillatory patterns during each iteration of the optimization loop. However, exerting precise control over a multitude of parameters within a high-dimensional latent space is an intricate challenge.
@@
Here, the generator function $G$ assumes an input latent vector $\bm{z}$ from the high-dimensional latent space $\bm{Z}$, and the evaluation function $E$ quantifies the quality of the generated output. In our specific context, the choice of the function $E$ is contingent upon the subjective perception of the user. The objective of the optimization is to discover the optimal latent vector $\bm{z}^{*}$ that maximizes the value of the evaluation function $E$ through human-guided exploration within the latent space [10].
@@ III-B Generator Design
We designed our vibrotactile generator architecture based on the framework of the super-resolution residual network (SRResNet) [24, 25]. Our motivation was to effectively address high-frequency image details. There is an input convolutional layer, 16 residual blocks, a mid convolutional layer, a 4-times upscale module consisting of two convolutional layers and two pixel shuffle layer, and an output convolutional layer in the SRResNet. We introduced a Fully Connected (FC) layer to the input end of the network as the resize layer for dimension transformation, to ensure that our latent vector can be received by SRResNet. Additionally, we made an adjustment to the final convolutional layer of the network to tailor the output to the required size, specifically to generate the desired spectrogram.
@@
To ensure the effectiveness of the generator, it is essential to employ an appropriate discriminator. A discriminator that performs exceptionally well or poor can hinder the training of the generator. In our study, we utilized a discriminator with a structure similar to the one presented in [24]. This discriminator comprises eight convolutional layers, albeit with adjustments to the channel numbers to align it with the specific requirements of our task. The comprehensive architecture of the GAN is shown in Fig. 3.
The generator and discriminator were trained simultaneously. In the training process, the generator $G$ learns how to generate new samples closer to the real sample, while the discriminator $D$ will discriminate the generated samples. The goal of the training can be represented as follows:
@@
where $z$ represents a random noise from the distribution $p(z)$ (Gaussian distribution in this research), and the $p_{data}(x)$ denotes the distribution of the real data $x$. This competitive process encourages the generator to generate samples that are indistinguishable from real data.
@@ III-C Generation with Conditional Knowledge
While GAN has outstanding generation performance, the training process is unstable. Additionally, GAN generate samples from random noise, resulting in uncontrollable outcomes. However, our objective is to exert control over the latent space, enabling the model to generate spectrograms that closely resemble real samples belonging to specific classes. To address these challenges, we built our model combining the advantages of GANs and autoencoders, and utilized conditional information. The inspiration for our model draws from previous works, such as Auxiliary Classifier GAN (ACGAN) [26], Autoencoding GAN (AEGAN) [27], and Conditional Adversarial Autoencoder (CAAE) [28] to enhance the training process and enable the model to generate samples belonging to particular classes.
The model shown in Fig. 2 contains four parts: encoder, generator, spectrogram discriminator, and latent discriminator. The encoder extracts the feature vector from the original spectrogram, while the generator utilizes this feature vector to reconstruct the spectrogram. Similar to the Adversarial Autoencoder (AAE) paradigm, a latent discriminator is employed to encourage the encoder’s output to conform to a predetermined prior distribution. This ensures that the latent space is evenly distributed according to the specified prior distribution, allowing the generator to effectively learn a mapping from this prior distribution to the spectrogram distribution, as proposed in [29]. Additionally, a determined distribution will benefit the following optimization of DSS. To make the feature containing richer information, we utilized the spectrogram discriminator as a classifier and applied the auxiliary classification loss, such as the ACGAN. It will work as the discriminator and classifier when receiving the encoded feature, and only as the discriminator when receiving the real gaussian noise.
To avoid data loss in compression and over-smoothing reconstruction, we used a generator together with the decoder loss, and use a weight value to balance them. The generator and spectrogram discriminator were structured according to the principles of CAAE. Notably, with the label as the conditional information, the generator can generate samples belonging to certain classes. The discriminator can learn the relationship in the pairs of the spectrogram and the corresponding class; thus, help the generator to provide correct spectrograms. The training goal can be described as follows:
@@
where $D_{z}$ represents the latent discriminator, $E$ and $G$ represent the encoder and generator, $D_{s}$ represents the spectrogram discriminator, and $C$ represent the auxiliary latent classifier, respectively. $x$ and $y$ represent the real data and the label. $TV$ represents the total variation loss. $\lambda$ and $\gamma$ are the weights for the decoder loss and the total variation loss, which can balance the smoothness and high resolution [28]. In this study, the weight setting is $\lambda=100$ and $\gamma=10$.
We initialized the encoder using a pre-trained ResNet-50 model that was trained on the ImageNet1K_V2 dataset, and introduced an FC layer to resize the spectrogram to the shape corresponding to the pre-trained model input. The generator and spectrogram discriminator follow the network structures illustrated in Fig. 3, maintaining consistency with those architectures. The latent discriminator contains two FC layers: 1 output layer for the discrimination and 1 output layer for the classification. To apply the label as conditional information, we used one-hot labels and used an FC layer to resize the label vector to the dimension corresponding to the feature map. The label information will be added to the feature map as a channel.